A salt contains only strontium and one of the halide ions. A 0.0910-g sample of the salt was dissolved in water, and an excess of sulfuric acid was added to form strontium sulfate (SrSO4), which was filtered, dried, and weighed. Its mass was found to be 0.133 g. What is the formula of the strontium halide? (Omit states-of-matter from your answer.)
Answer #1

The balanced equation between SrX2 and H2SO4 can be written as follows.

SrX2 + H2SO4 \rightarrow SrSO4 + 2HX

Here, the no. of moles of SrSO4 formed = 0.091 g/(183.676 g/mol) = 7.241×10^-4 mol

According to the equation shown above, the mole ratio of SrX2: SrSO4 = 1: 1

Therefore, the no. moles of SrX2 = 7.241×10^-4 mol

Hence, the molar mass of SrX2 = 0.091 g/(7.241×10^-4 mol) = 125.6 g/mol = molar mass of SrF2

i.e. The formula of strontium halide is SrF2